### 01 Apr 2026 v1

# OmniMem: Autoresearch-Guided Discovery of Lifelong Multimodal Agent Memory

Abstract
AI agents increasingly operate over extended time horizons, yet their ability to retain, organize, and recall multimodal experiences remains a critical bottleneck. Building effective lifelong memory requires navigating a vast design space spanning architecture, retrieval strategies, prompt engineering, and data pipelines; this space is too large and interconnected for manual exploration or traditional AutoML to explore effectively. We deploy an autonomous research pipeline to discover OmniMem, a unified multimodal memory framework for lifelong AI agents. Starting from a naïve baseline (F1 = 0.117 on LoCoMo), the pipeline autonomously executes experiments across two benchmarks, diagnosing failure modes, proposing architectural modifications, and repairing data pipeline bugs, all without human intervention in the inner loop. The resulting system achieves state-of-the-art on both benchmarks, improving F1 by +411% on LoCoMo (0.1170.598) and +214% on Mem-Gallery (0.2540.797) relative to the initial configurations. Critically, the most impactful discoveries are not hyperparameter adjustments: bug fixes (+175%), architectural changes (+44%), and prompt engineering (+188% on specific categories) each individually exceed the cumulative contribution of all hyperparameter tuning, demonstrating capabilities fundamentally beyond the reach of traditional AutoML. We provide a taxonomy of six discovery types and identify four properties that make multimodal memory particularly suited for autoresearch, offering guidance for applying autonomous research pipelines to other AI system domains. Code is available at this https://github.com/aiming-lab/OmniMem.
1 Introduction
Recent advances in large language models have given rise to AI agents capable of tool use, multi-step reasoning, and cross-modal comprehension (Yao et al., 2023; Yan et al., 2025). These agents interact with users over extended time horizons, accumulating diverse streams of text, images, audio, and video throughout their operation. However, their ability to retain, organize, and recall past experiences remains a critical bottleneck (Zhang et al., 2024; Xu et al., 2025b). Building effective lifelong multimodal memory requires navigating a vast design space spanning architectural choices (how to structure storage), retrieval strategies (how to find relevant information), prompt engineering (how to present context to the LLM), and data pipeline configurations (how to ingest and process heterogeneous inputs).
Existing approaches to agent memory fall into two broad categories, each with notable limitations. The first stores raw inputs and retrieves them via embedding similarity (Lewis et al., 2020; Borgeaud et al., 2022), suffering from storage bloat and retrieval noise as the memory grows. The second introduces structured memory management with explicit operations (Packer et al., 2023; Park et al., 2023), but typically operates on text alone, discarding rich visual and auditory signals. Crucially, both categories are products of manual research cycles: a human researcher hypothesizes an improvement, implements it, evaluates on a benchmark, and iterates. A single researcher may explore only a handful of configurations per day, and important interactions between tightly coupled components are easily missed. Traditional AutoML methods (Hutter et al., 2019a) can search over predefined numerical hyperparameter spaces, but cannot perform the code comprehension, bug diagnosis, architectural redesign, and cross-component reasoning that account for the largest performance gains in complex systems. As a result, existing memory systems inherit the blind spots of their designers–limitations that a more systematic search could avoid.
Recent work on autonomous scientific discovery (Lu et al., 2024; Romera-Paredes et al., 2024; Panfilov et al., 2026) has shown that LLM agents can autonomously discover novel algorithms that outperform human-designed baselines, provided the target domain admits well-defined, quantitative evaluation signals. We ask whether this paradigm extends to complex, multi-component AI systems and answer affirmatively. We deploy AutoResearchClaw (Liu et al., 2026b), a 23-stage autonomous research pipeline, to discover OmniMem, a unified multimodal memory framework for lifelong AI agents. Starting from a naïve baseline (F1 = 0.117 on LoCoMo), the pipeline autonomously executes experiments across two benchmarks, iteratively diagnosing failure modes, proposing architectural modifications, repairing data pipeline bugs, and validating improvements, all without human intervention in the inner loop. The resulting system achieves state-of-the-art on both benchmarks, improving F1 by +411% on LoCoMo (0.1170.598) and +214% on Mem-Gallery (0.2540.797) relative to the initial configurations. Critically, the most impactful discoveries are not hyperparameter adjustments: bug fixes (+175%), architectural changes (+44%), and prompt engineering (+188% on specific categories) each individually exceed the cumulative contribution of all hyperparameter tuning, demonstrating capabilities fundamentally beyond the reach of traditional AutoML.
Among the pipeline’s most consequential discoveries are three architectural principles that define OmniMem. First, selective ingestion: lightweight perceptual encoders measure the information novelty of each incoming signal and discard redundant content before storage, significantly reducing storage requirements. Second, unified representation: all memories, regardless of modality, are represented as Multimodal Atomic Units (MAUs) that separate lightweight metadata from heavy raw data, enabling fast search over compact metadata while preserving full-content access on demand. Third, progressive retrieval: a pyramid mechanism expands information in three stages (summaries, details, raw evidence), each gated by a token budget, backed by a hybrid search strategy combining dense vector retrieval with sparse keyword matching via set-union merging, a strategy autonomously discovered by the pipeline. Our key observation is that multimodal memory is particularly well-suited for autonomous research pipelines due to four properties: immediate scalar evaluation metrics enabling tight optimization loops, modular architecture allowing isolated component modification, fast iteration cycles (1–2 hours per experiment) supporting dozens of hypotheses within days, and version-controlled code modifications allowing failed experiments to be cleanly reverted.
In summary, our primary contribution is OmniMem, a unified multimodal memory framework whose architecture and configuration are discovered through AutoResearchClaw and that achieves state-of-the-art results on both evaluated benchmarks. Beyond the system itself, we provide a comprehensive taxonomy of autonomous discoveries across experiments, revealing that the highest-impact improvements lie beyond the reach of traditional AutoML, and we characterize the convergence behavior, failure modes, and automatic recovery patterns of the pipeline. Our analysis further identifies four properties that make multimodal memory a particularly suitable domain for autoresearch, providing guidance for future applications to other AI system domains.
2 Related Work
Autonomous Scientific Discovery. The vision of AI-driven research has advanced rapidly. The AI Scientist (Lu et al., 2024) demonstrated end-to-end paper generation at per paper across three ML domains, with its successor AI Scientist v2 (Yamada et al., 2025) eliminating human-authored templates through agentic tree search. FunSearch (Romera-Paredes et al., 2024) pairs LLM creativity with programmatic evaluation to discover novel mathematical constructions. AutoResearchClaw (Liu et al., 2026b) introduces a 23-stage autonomous research pipeline with multi-agent debate and self-healing execution. AI-Researcher (Tang et al., 2025) introduces collaborative multi-agent frameworks with structured iterative refinement, and Bilevel Autoresearch (Qu and Lu, 2026) meta-optimizes the search strategy itself. A comprehensive survey by Tie et al. (2026) maps the evolution from foundational modules (2022–2023) through closed-loop systems (2024) to scalable human-AI collaboration (2025+). We apply the autoresearch paradigm to multi-component AI system optimization, where the challenge shifts from discovering isolated artifacts to diagnosing and improving interactions across tightly coupled modules.
Multimodal Memory Systems. Memory-augmented LLM agents have evolved from text-only systems, including MemGPT (Packer et al., 2023) with OS-inspired memory hierarchies, Generative Agents (Park et al., 2023) with recency-importance-relevance scoring, SimpleMem (Liu et al., 2026a) with efficient lifelong memory, and A-Mem (Xu et al., 2025a) with LLM-directed reorganization, to multimodal architectures. MemVerse (Liu et al., 2025) combines episodic-semantic memory with multimodal knowledge graphs but requires three LLM calls per ingested item. Mem0 (Chhikara et al., 2025) offers dynamic fact extraction with optional graph memory. VisRAG (Yu et al., 2025) indexes visual pages directly, avoiding text extraction losses. Claude-Mem (Anthropic, 2024) provides commercial embedding-based dialogue memory. These systems all require extensive manual tuning of retrieval strategies, ingestion pipelines, and prompt configurations, which is precisely the kind of optimization that autonomous research pipelines can accelerate.
Automated Machine Learning. Neural Architecture Search (Hutter et al., 2019b; Zoph and Le, 2017; Liu et al., 2019) automates model design but operates primarily on well-defined architectural search spaces with differentiable or reinforcement-learning-based objectives. Hyperparameter optimization methods (Bergstra et al., 2011; Falkner et al., 2018) efficiently navigate continuous and categorical spaces, while systems like Auto-sklearn 2.0 (Feurer et al., 2022) automate full ML pipelines including preprocessing and model selection via meta-learning. More recently, LLM-based agents have been applied to ML tasks: MLAgentBench (Huang et al., 2024) benchmarks LLM agents on ML research tasks involving code modification, demonstrating the potential of language-guided optimization. Our setting differs fundamentally: the “search space” includes not only hyperparameters and architectural choices, but also prompt engineering, data pipeline bug detection and repair, evaluation format alignment, and cross-component interaction diagnosis, all of which require natural language understanding and code modification capabilities beyond traditional AutoML.
3 Autoresearch-Guided Discovery of OmniMem
In this section, we describe the autonomous optimization process and the system it produces. We first overview the pipeline (§3.1), then present the discovered OmniMem architecture (§3.2), followed by the benchmark-specific optimization strategy (§3.3).
3.1 Pipeline Overview
As discussed in Section 1, the design space of multimodal memory systems is too large and interconnected for manual exploration to cover effectively. To address this, we deploy AutoResearchClaw (Liu et al., 2026b), a 23-stage autonomous research pipeline, to systematically optimize OmniMem. The pipeline receives three inputs: (1) the SimpleMem (Liu et al., 2026a) codebase, a unimodal text-only lifelong memory framework, as a starting point, (2) two benchmark evaluation harnesses with quantitative metrics (F1), and (3) API access to LLM providers. It then enters an iterative loop: at each step, the pipeline analyzes prior results, generates a hypothesis for improvement, implements the change in code, evaluates on a benchmark, and decides whether to proceed (metric improved by ), iterate (ambiguous result; refine the current hypothesis), or pivot (two consecutive degradations; revert and try a new direction). Of total experiments, most resulted in a proceed decision, with the remainder split between iterate and pivot. The full pipeline phases (scoping, literature discovery, multi-agent debate, experiment design, sandboxed execution, analysis, documentation, and finalization) are described in Liu et al. (2026b); here we focus on the discoveries it makes and the mechanisms that enable them.
3.2 The Discovered Architecture
The pipeline takes SimpleMem (Liu et al., 2026a), a unimodal text-only lifelong memory framework, as its starting point. We provide AutoResearchClaw with the SimpleMem codebase and instruct it to extend the system from text-only memory to full multimodal support, autonomously designing the necessary architectural components for ingesting, storing, and retrieving heterogeneous signals (text, images, audio, video). Through iterative experimentation, the pipeline converges to an architecture organized around three principles: selective ingestion, progressive retrieval, and structured knowledge (Figure 2).
3.2.1 Selective Ingestion
The first principle is selective ingestion: the system first filters redundant inputs, then encapsulates the retained signals into a unified multimodal representation.
Novelty-Based Filtering. Before any data enters the memory store, lightweight perceptual encoders assess the novelty of incoming information and discard redundant content. For vision, CLIP embeddings are compared across consecutive frames to detect scene changes; for audio, VAD speech probability gates retention to reject silence; for text, Jaccard overlap with recent summaries filters near-duplicates. This filtering significantly reduces storage requirements without losing semantic content.
Multimodal Atomic Units. Signals that pass the novelty filter are encapsulated as Multimodal Atomic Units (MAUs), , which decouple compact searchable metadata from heavy raw content. Here is a text summary, is its embedding, points to raw content in cold storage, is the timestamp, is modality, and stores structural links to other MAUs. This yields a two-tier design: hot storage keeps summaries, embeddings, and temporal/graph metadata for fast retrieval, while cold storage keeps large assets (images, audio, video) and is accessed lazily through .
3.2.2 Progressive Retrieval with Hybrid Search
Once memories are ingested and stored as MAUs, the next challenge is how to retrieve them efficiently at query time. The second principle is progressive retrieval: rather than loading all retrieved content into the LLM context at once, OmniMem expands information in stages under explicit token budgets.
Hybrid Dense-Sparse Search. Given a user query , dense retrieval via FAISS (Johnson et al., 2021), a library for efficient similarity search over high-dimensional vectors, yields semantically similar candidates by inner-product search over L2-normalized MAU embeddings. In parallel, BM25 (Robertson and Zaragoza, 2009) scoring over MAU summaries yields keyword-matched candidates . A key discovery of the autonomous pipeline is set-union merging: empirically, score-based re-ranking (the standard approach) disrupts semantic ordering and degrades performance. Instead, dense results retain their original ranking and BM25-only results are appended:
Pyramid Retrieval. The hybrid search above produces a candidate set , each candidate scored once by cosine similarity during dense retrieval. The pyramid mechanism then progressively expands the content of these candidates across three levels, reusing this score to gate each transition: Level 1 returns only summaries (10 tokens each) for the top- candidates by similarity; Level 2 loads full text or detailed captions for candidates whose similarity exceeds a threshold ; Level 3 loads raw content (images, audio) from cold storage under an explicit token budget B, expanding items greedily in decreasing similarity-per-token order. All transitions are governed by deterministic rules rather than LLM judgment, avoiding additional latency while adapting context depth to each query’s complexity.
3.2.3 Knowledge Graph-Augmented Retrieval
While hybrid search and pyramid retrieval handle queries that can be answered from individual MAUs, many real-world queries require reasoning over multiple connected facts (e.g., “What gift did I give to the person I met at the conference in March?”). The third principle is therefore structured knowledge: OmniMem maintains a knowledge graph that captures entities and relationships across all MAUs.
During MAU creation, an LLM extracts entities and directed relations from each summary, producing entity-relation triples. Each entity carries a type label from 7 categories (Person, Location, Event, Concept, Time, Organization, Object) and is linked back to its source MAU. As new MAUs are ingested, the same real-world entity may appear under different surface forms (e.g., “Dr. Smith” vs. “John Smith”). To prevent node fragmentation, entity resolution merges entities whose hybrid similarity, combining cosine similarity over name embeddings with Jaro-Winkler string similarity, exceeds a threshold.
At query time, the system identifies seed entities mentioned in the query and performs bounded neighborhood expansion within hops. Each reached entity is scored with distance-decayed relevance , where is the shortest path distance to any seed entity and is a decay factor. MAUs linked to high-scoring graph entities are merged with the hybrid search results from , providing both direct content matches and relationally connected evidence for answer generation.
3.3 Benchmark-Specific Optimization
Having described the multimodal architecture that the pipeline discovers from SimpleMem, we now turn to how it optimizes this architecture for each target benchmark. The pipeline employs a two-phase strategy: rapid iteration on a small training subset, followed by evaluation on the held-out test set.
Development subset for fast iteration. For each benchmark, the pipeline selects a small representative subset for rapid experimentation during the optimization loop. On LoCoMo, a small subset of conversations is used for iterative development, enabling each experiment to complete in under 2 hours. On Mem-Gallery, a small subset of datasets is used, with each experiment completing in minutes. This design enables the pipeline to explore dozens of hypotheses within days. After the optimization trajectory converges, the final configuration is evaluated on the complete benchmark to ensure generalization and to maintain consistency with the evaluation protocols used by prior memory systems.
Iterative diagnosis and repair. During each optimization cycle, the pipeline autonomously diagnoses and repairs failures at two levels. At the execution level, when an experiment fails or produces unexpected outputs, a self-healing module classifies the error (API error, dependency error, runtime exception, output format mismatch) and generates a targeted fix. For example, when the embedding service returned 403 errors due to an expired API key, the module detected the authentication failure pattern and switched to a local sentence-transformer backend without manual intervention. At the semantic level, when experiments succeed but produce unexpectedly poor metrics, the pipeline performs deeper analysis.
4 Experiments
We evaluate OmniMem along two dimensions: (1) the autonomous optimization process, specifically whether the pipeline discovers meaningful improvements across diverse benchmarks, and (2) final system quality, examining whether the discovered architecture achieves state-of-the-art results and whether individual components contribute meaningfully.
4.1 Experimental Setup
Benchmarks. We evaluate on two benchmarks spanning complementary types of memory-dependent reasoning: LoCoMo (Maharana et al., 2024) (1,986 QA pairs across multi-session dialogues, token-level F1) and Mem-Gallery (Bei et al., 2026) (1,711 QA pairs from 240 multimodal dialogues with 1,003 grounded images, F1). Detailed benchmark descriptions are provided in Appendix A.
Baselines. We compare against six memory systems representing diverse design philosophies: MemVerse (Liu et al., 2025) (hierarchical episodic-semantic memory with multimodal knowledge graph), Mem0 (Chhikara et al., 2025) (dynamic fact extraction with optional graph memory), Claude-Mem (Anthropic, 2024) (commercial embedding-based dialogue memory), A-MEM (Xu et al., 2025a) (LLM-directed memory reorganization), MemGPT (Packer et al., 2023) (OS-inspired memory hierarchies), and SimpleMem (Liu et al., 2026a) (efficient lifelong memory with atomization and adaptive pruning). All are evaluated under identical splits and protocols.
Implementation. Dense retrieval uses FAISS with all-MiniLM-L6-v2 embeddings (384d); sparse retrieval uses BM25; visual novelty filtering uses frozen CLIP ViT-B/32. Knowledge graph extraction uses GPT-4o in JSON mode. Default configuration: top-=20, =0.4, =6,000 tokens. Full per-benchmark configuration details are provided in Appendix C.
4.2 Optimization Trajectories
| LoCoMo | Mem-Gallery | |||||||||||
| Backbone | Method | MH | SH | Tmp | Open | Adv | All | F1 | EM | B | B-1 | B-2 |
| GPT-4o | MemVerse | 0.260 | 0.157 | 0.196 | 0.192 | 0.944 | 0.365 | 0.505 | 0.330 | 0.270 | 0.440 | 0.355 |
| Mem0 | 0.309 | 0.156 | 0.217 | 0.295 | 0.857 | 0.397 | 0.298 | 0.192 | 0.182 | 0.268 | 0.224 | |
| Claude-Mem | 0.294 | 0.153 | 0.167 | 0.243 | 0.915 | 0.383 | 0.210 | 0.148 | 0.148 | 0.194 | 0.170 | |
| A-MEM | 0.295 | 0.174 | 0.200 | 0.266 | 0.898 | 0.394 | 0.370 | 0.252 | 0.240 | 0.332 | 0.285 | |
| MemGPT | 0.305 | 0.188 | 0.246 | 0.305 | 0.843 | 0.404 | 0.435 | 0.298 | 0.275 | 0.390 | 0.335 | |
| SimpleMem | 0.318 | 0.195 | 0.235 | 0.308 | 0.802 | 0.432 | 0.535 | 0.348 | 0.310 | 0.468 | 0.390 | |
| OmniMem | 0.556 | 0.365 | 0.255 | 0.641 | 0.835 | 0.598 | 0.797 | 0.449 | 0.366 | 0.627 | 0.505 | |
| GPT-4o-mini | MemVerse | 0.147 | 0.074 | 0.106 | 0.093 | 0.747 | 0.290 | 0.450 | 0.295 | 0.248 | 0.395 | 0.330 |
| Mem0 | 0.285 | 0.112 | 0.179 | 0.297 | 0.761 | 0.364 | 0.291 | 0.188 | 0.185 | 0.265 | 0.223 | |
| Claude-Mem | 0.245 | 0.102 | 0.122 | 0.215 | 0.845 | 0.338 | 0.272 | 0.175 | 0.172 | 0.245 | 0.210 | |
| A-MEM | 0.278 | 0.091 | 0.163 | 0.260 | 0.823 | 0.357 | 0.330 | 0.222 | 0.205 | 0.298 | 0.252 | |
| MemGPT | 0.283 | 0.113 | 0.182 | 0.289 | 0.776 | 0.364 | 0.398 | 0.262 | 0.242 | 0.355 | 0.298 | |
| SimpleMem | 0.300 | 0.128 | 0.178 | 0.312 | 0.891 | 0.404 | 0.498 | 0.318 | 0.290 | 0.435 | 0.368 | |
| OmniMem | 0.544 | 0.196 | 0.177 | 0.588 | 0.779 | 0.519 | 0.749 | 0.403 | 0.334 | 0.583 | 0.465 | |
| GPT-4.1-nano | MemVerse | 0.146 | 0.061 | 0.169 | 0.115 | 0.711 | 0.256 | 0.470 | 0.308 | 0.255 | 0.410 | 0.340 |
| Mem0 | 0.290 | 0.134 | 0.194 | 0.277 | 0.537 | 0.310 | 0.268 | 0.176 | 0.156 | 0.238 | 0.199 | |
| Claude-Mem | 0.087 | 0.029 | 0.119 | 0.047 | 0.705 | 0.246 | 0.303 | 0.194 | 0.172 | 0.268 | 0.223 | |
| A-MEM | 0.045 | 0.016 | 0.142 | 0.050 | 0.747 | 0.216 | 0.365 | 0.242 | 0.225 | 0.325 | 0.275 | |
| MemGPT | 0.287 | 0.130 | 0.234 | 0.279 | 0.556 | 0.316 | 0.360 | 0.238 | 0.218 | 0.318 | 0.268 | |
| SimpleMem | 0.298 | 0.145 | 0.210 | 0.285 | 0.648 | 0.342 | 0.518 | 0.338 | 0.300 | 0.452 | 0.380 | |
| OmniMem | 0.477 | 0.216 | 0.244 | 0.583 | 0.722 | 0.492 | 0.780 | 0.430 | 0.353 | 0.610 | 0.488 | |
| GPT-5.1 | MemVerse | 0.287 | 0.173 | 0.277 | 0.297 | 0.780 | 0.383 | 0.478 | 0.312 | 0.262 | 0.418 | 0.345 |
| Mem0 | 0.292 | 0.160 | 0.261 | 0.298 | 0.819 | 0.390 | 0.270 | 0.175 | 0.157 | 0.240 | 0.200 | |
| Claude-Mem | 0.289 | 0.171 | 0.264 | 0.292 | 0.814 | 0.388 | 0.305 | 0.203 | 0.188 | 0.279 | 0.230 | |
| A-MEM | 0.287 | 0.164 | 0.246 | 0.284 | 0.826 | 0.385 | 0.408 | 0.268 | 0.242 | 0.365 | 0.302 | |
| MemGPT | 0.288 | 0.165 | 0.249 | 0.294 | 0.806 | 0.385 | 0.425 | 0.275 | 0.250 | 0.378 | 0.315 | |
| SimpleMem | 0.305 | 0.178 | 0.272 | 0.305 | 0.807 | 0.418 | 0.538 | 0.350 | 0.312 | 0.470 | 0.395 | |
| OmniMem | 0.598 | 0.367 | 0.307 | 0.676 | 0.747 | 0.613 | 0.810 | 0.460 | 0.374 | 0.639 | 0.515 | |
| GPT-5-nano | MemVerse | 0.208 | 0.203 | 0.168 | 0.252 | 0.741 | 0.366 | 0.478 | 0.315 | 0.262 | 0.420 | 0.345 |
| Mem0 | 0.264 | 0.143 | 0.237 | 0.270 | 0.737 | 0.352 | 0.283 | 0.176 | 0.165 | 0.250 | 0.210 | |
| Claude-Mem | 0.091 | 0.063 | 0.092 | 0.088 | 0.736 | 0.275 | 0.350 | 0.249 | 0.217 | 0.315 | 0.264 | |
| A-MEM | 0.260 | 0.149 | 0.223 | 0.257 | 0.745 | 0.348 | 0.505 | 0.332 | 0.290 | 0.445 | 0.368 | |
| MemGPT | 0.267 | 0.151 | 0.226 | 0.271 | 0.744 | 0.355 | 0.388 | 0.255 | 0.230 | 0.345 | 0.288 | |
| SimpleMem | 0.278 | 0.200 | 0.245 | 0.282 | 0.824 | 0.388 | 0.522 | 0.340 | 0.302 | 0.458 | 0.385 | |
| OmniMem | 0.357 | 0.371 | 0.253 | 0.561 | 0.719 | 0.522 | 0.787 | 0.437 | 0.357 | 0.617 | 0.494 | |
Figure 3 visualizes the optimization trajectories. The pipeline completed experiments across both benchmarks in 72 hours of wall-clock time, a coverage that would require approximately 4 weeks for a human researcher at 3 experiments per day. We highlight the most impactful discoveries; detailed per-iteration tables are provided in Appendix F.
LoCoMo (9 iterations, F1: 0.1170.598). The pipeline executed 9 successful iterations over 48 hours, with 2 additional experiments automatically reverted (full trajectory in Appendix Table 6). The most impactful discovery (Iter 1, +175%) was identifying that the API call lacked a response_format parameter, a one-line bug causing 9 verbosity that destroyed F1 precision. In Iter 5, the pipeline discovered that all 4,277 MAU timestamps had been corrupted to the ingestion date and autonomously generated a keyword-matching script that corrected 99.98% of them without re-ingestion. The pipeline also discovered that set-union merging of FAISS and BM25 results (Iter 2) significantly outperforms score-based fusion, a finding confirmed by ablation (Section 4.4.1).
Mem-Gallery (39 experiments, F1: 0.2540.797). Optimization spanned 7 phases (full trajectory in Appendix Table 7). The single largest improvement (+53%) came from discovering that returning full original dialogue text instead of LLM-generated summaries dramatically improves token-overlap F1, a non-obvious finding, since summaries are traditionally preferred for efficiency. The pipeline also found that prompt constraint positioning (before vs. after the question) matters more than constraint content, with one category improving +188% from this change alone. After Phase 7, four independent runs yielded F1 in [0.791, 0.797], confirming the performance ceiling and triggering the pipeline’s decision to stop.
4.3 Main Results
To contextualize these results against existing memory systems, we conduct a controlled comparison of OmniMem against six baselines across five LLM backbones (GPT-4o, GPT-4o-mini, GPT-4.1-nano, GPT-5.1, and GPT-5-nano). Table 1 reports per-category F1 on LoCoMo and five evaluation metrics on Mem-Gallery across all backbones.
On LoCoMo, OmniMem achieves the highest overall F1 across all backbones, ranging from 0.492 (GPT-4.1-nano) to 0.613 (GPT-5.1), substantially outperforming SimpleMem (0.342–0.432), the current state-of-the-art on LoCoMo (Liu et al., 2026a). OmniMem dominates on multi-hop, single-hop, and open-domain categories, with particularly large margins on open-domain questions.
On Mem-Gallery, OmniMem achieves F1 ranging from 0.749 to 0.810, consistently outperforming all memory baselines by a wide margin. SimpleMem is again the strongest baseline (F1 up to 0.538 with GPT-5.1), but still trails OmniMem by over 25 percentage points. These patterns confirm that OmniMem’s gains come from its architectural design (hybrid search, pyramid retrieval, knowledge graph augmentation) rather than from a single dominant component.
4.4 Analysis
4.4.1 Ablation Studies
| Component Removed | F1 | Rel. |
| w/o Pyramid Expansion | 10.2 | 17% |
| w/o BM25 Hybrid | 8.5 | 14% |
| w/o LLM Summarization | 7.3 | 12% |
| Reduced top- (5 vs 20) | 4.2 | 7% |
| w/o Metadata Context | 1.4 | 2% |
Table 2 presents an ablation study on LoCoMo that validates key design choices discovered by the pipeline. Specifically, we remove individual components and report the mean F1 across 4 LLM backbones. Pyramid expansion (17%) and BM25 hybrid search (14%) are the most critical, and both were significantly refined by the autonomous pipeline during optimization. LLM summarization contributes 12%, confirming that compact MAU summaries are essential for retrieval quality. Reducing top- from 20 to 5 costs 7%, while metadata context has a modest effect (2%). Notably, the two most impactful components (pyramid expansion and hybrid search) are precisely those that received the most optimization iterations, suggesting the pipeline correctly allocated its search budget.
4.4.2 Efficiency
OmniMem achieves 5.81 queries/sec with 8 parallel workers (3.5 faster than the fastest baseline), enabled by read-only FAISS and BM25 indices supporting concurrent lookup (Figure 4, Table 3). All baselines are bottlenecked by sequential LLM generation (85–97% of per-query time), while OmniMem parallelizes the retrieval-generation pipeline via thread-safe read-only indices.
| Method | q/s | Ret. | Gen. |
| SimpleMem | 1.68 | 45 | 550 |
| MemVerse | 1.48 | 70 | 596 |
| Mem0 | 1.46 | 18 | 665 |
| OmniMem (w=1) | 1.05 | 118 | 846 |
| OmniMem (w=8) | 5.81 | 461 | 821 |
| Ret./Gen. in milliseconds. | |||
4.4.3 Case Study: Multi-Hop Retrieval
We illustrate OmniMem’s retrieval pipeline on a real multi-hop query from LoCoMo that requires synthesizing facts across separate conversation sessions. The query asks: “What subject have Caroline and Melanie both painted?” The gold answer is “sunsets,” but answering correctly requires retrieving each person’s painting history from different sessions and identifying the overlap.
Hybrid search. Dense retrieval returns MAUs mentioning Caroline’s paintings (e.g., “Caroline painted a sunset scene”) and Melanie’s art projects (e.g., “Mel and her kids painted a sunset with a palm tree”), but these appear in separate sessions with different surrounding context. BM25 recovers additional MAUs containing the keyword “paint” that rank lower in dense results. Set-union merging preserves the dense ordering and appends BM25-only matches.
Knowledge graph expansion. The query processor extracts seed entities Caroline (Person) and Melanie (Person). Neighborhood expansion links both entities to painting (Concept) and sunset (Concept) through separate relation paths, surfacing MAUs that mention each person’s painting activities even when the surface text does not co-mention both names.
Pyramid retrieval and answer. Level 1 summaries from both relation paths are loaded; their similarity scores exceed , triggering Level 2 expansion of the full conversation text. The LLM identifies “sunsets” as the common subject and produces the correct answer (F1 = 1.0). In contrast, MemGPT, which lacks cross-session entity linking, hallucinates “Horses” (F1 = 0.0).
5 Conclusion
We have presented OmniMem, a unified multimodal memory framework whose architecture and configuration are discovered through AutoResearchClaw, an autonomous research pipeline. Starting from a naïve baseline (F1 = 0.117 on LoCoMo), the pipeline autonomously executed experiments in 72 hours, achieving state-of-the-art on both LoCoMo and Mem-Gallery. The highest-impact discoveries, including bug fixes (+175%), architectural changes (+44%), and prompt engineering (+188% on specific categories), require code comprehension and cross-component reasoning beyond the reach of traditional AutoML. Our taxonomy of six discovery types, together with the observation that multimodal memory is well-suited for autoresearch due to its scalar metrics, modular architecture, and fast iteration cycles, provides a roadmap for applying autonomous research pipelines to other complex AI system domains.
Ethics Statement
OmniMem persistently retains multimodal personal information (text, images, audio, video) over extended time horizons, raising important considerations regarding data privacy and consent. Real-world deployment necessitates explicit user consent, transparent retention policies, and robust mechanisms to review, modify, or delete stored memories. The knowledge graph component, which extracts and links entities across memory units, could enable profiling if deployed without adequate access controls and encryption. Regarding the autonomous research pipeline, all AutoResearchClaw experiments are conducted on established academic benchmarks with no real user data; deploying such pipelines on production systems would require additional oversight. We advocate for memory-augmented agents that prioritize user autonomy, data minimization, and the right to be forgotten.
Reproducibility Statement
We provide complete hyperparameters (Table 4), per-benchmark configurations (Table 5), and system architecture details in Appendix C. Full optimization trajectories with per-iteration metrics and decisions are documented in Appendix E. All prompts are reproduced verbatim in Appendix G, and the AutoResearchClaw pipeline stages are described in Appendix D. Both benchmarks, LoCoMo (Maharana et al., 2024) and Mem-Gallery (Bei et al., 2026), are publicly available; we evaluate under identical splits and protocols as prior work using standard metrics (token-level F1, exact match, BLEU). We open-source the complete framework, benchmark harnesses, and full experiment logs to enable independent replication.
Appendix A Benchmark Details
LoCoMo.
The Long-term Conversation Memory benchmark (Maharana et al., 2024) evaluates an agent’s ability to recall and reason over extended multi-session dialogues. It comprises 10 conversations with 19–32 sessions each (K tokens per dialogue), generating 1,986 QA pairs across five categories: Single-hop (SH) requires retrieving a single fact; Multi-hop (MH) requires synthesizing information across multiple sessions; Temporal (T) tests reasoning about when events occurred; Open-ended (O) requires generating longer, contextual responses; and Adversarial (A) tests the ability to correctly refuse unanswerable questions. Evaluation uses token-level F1 with Porter stemming.
Mem-Gallery.
Mem-Gallery (Bei et al., 2026) evaluates multimodal long-term memory in social interactions, comprising 1,711 QA pairs from 240 multi-session dialogues with 1,003 grounded images and 3,962 conversational rounds. Questions span 9 categories: Action Recognition (AR), Compound Decomposition (CD), Visual Search (VS), Timeline Learning (TTL), Temporal Reasoning (TR), Fact Retrieval (FR), Visual Reasoning (VR), Knowledge Reasoning (KR), and Multi-Entity Reasoning (MR). We use the public 20-dataset subset from HuggingFace.
Appendix B Baseline Descriptions
MemVerse
(Liu et al., 2025) combines episodic and semantic memory layers with a multimodal knowledge graph. Each ingested item requires 3 LLM calls (summarization, entity extraction, relation linking), yielding an ingestion rate of 0.22 items/sec. Retrieval uses embedding-based similarity over the knowledge graph.
Mem0
(Chhikara et al., 2025) performs dynamic fact extraction from conversational inputs, storing structured facts with optional graph memory augmentation. It achieves 1.28 items/sec ingestion and 18ms search latency, but is fundamentally text-only.
SimpleMem
(Liu et al., 2026a) introduces efficient lifelong memory management through three core mechanisms: memory atomization (decomposing inputs into fine-grained atomic units), adaptive consolidation (merging related memories to reduce redundancy), and context-aware pruning. It achieves state-of-the-art performance on LoCoMo with significantly reduced token cost (45% less than Mem0) and 4 faster total processing time.
Claude-Mem
(Anthropic, 2024) provides commercial embedding-based dialogue memory with per-turn image storage.
A-MEM
(Xu et al., 2025a) introduces agentic memory management where the LLM itself decides when and how to reorganize stored memories, including operations such as merging, splitting, and abstracting memory entries to maintain a compact and relevant memory state.
MemGPT
(Packer et al., 2023) draws an analogy to operating system memory hierarchies, implementing a main context (analogous to RAM) and external storage (analogous to disk) with explicit push/pop operations managed by the LLM agent. It achieves strong performance on conversational tasks but is fundamentally text-only and bottlenecked by sequential LLM calls for each memory management decision.
Appendix C OmniMem Technical Details
C.1 System Architecture
The complete OmniMem implementation comprises 13,300 lines of Python across 11 subpackages:
Core (MAU data structure, event hierarchy, unified configuration)
Processors (modality-specific ingestion: text, image, audio, video)
Storage (three-layer persistence: MAU store as JSON Lines, FAISS vector store, cold storage on filesystem/S3)
Retrieval (pyramid retriever, query processor, expansion manager)
Knowledge (entity extractor, in-memory knowledge graph, graph retriever)
Orchestrator (central coordinator, 873 lines, unified API)
C.2 Hyperparameters and Benchmark-Specific Configurations
Table 4 lists the complete hyperparameter specification. Table 5 summarizes the per-benchmark configurations discovered by the pipeline.
| Component | Parameter | Value |
| Visual Ingestion | CLIP threshold | 0.9 |
| CLIP threshold | 0.7 | |
| Frame buffer size | 3 | |
| Audio Ingestion | VAD threshold | 0.5 |
| Text Ingestion | Jaccard threshold | 0.8 |
| Entity Resolution | Merge threshold | 0.85 |
| Semantic weight | 0.5 | |
| Graph decay | 0.7 | |
| Expansion hops | 2 | |
| Pyramid Retrieval | Auto-expand threshold | 0.4 |
| Token budget | 6,000 | |
| BM25 parameters , | 1.5, 0.75 |
| Parameter | LoCoMo | Mem-Gallery |
| Embedding | text-emb-3-large | MiniLM-L6-v2 |
| Embedding dim | 3072 | 384 |
| top- | 20–30 | 20–40† |
| Token budget | 6,000 | 2K–8K† |
| Per-doc memory | No | Yes |
| System prompt | Yes | Yes |
C.3 Knowledge Graph Schema
The knowledge graph uses 7 entity types (Person, Location, Object, Event, Concept, Time, Organization) and 7 relation types (located_in, part_of, interacts_with, owns, attended, created_by, related_to). Entity extraction is performed by GPT-4o in JSON mode during MAU creation.
Appendix D AutoResearchClaw Pipeline Details
The complete pipeline comprises 23 stages organized in 8 phases:
A.
Research Scoping (Stages 1–2): Formulates SMART research goals; auto-detects hardware (GPU type, VRAM).
B.
Literature Discovery (Stages 3–6): Queries OpenAlex, Semantic Scholar, arXiv; screens by relevance/quality (gate at Stage 5); extracts structured knowledge cards.
C.
Knowledge Synthesis (Stages 7–8): Clusters findings; generates falsifiable hypotheses via multi-agent debate.
D.
Experiment Design (Stages 9–11): Designs protocols (gate at Stage 9); generates hardware-aware Python code with AST validation; schedules resources.
E.
Experiment Execution (Stages 12–13): Runs experiments in sandbox/Docker/SSH; self-healing loop (up to 10 retries) with structured deficiency classification.
F.
Analysis & Decision (Stages 14–15): Statistical analysis (t-tests, bootstrap CI); autonomous PROCEED/PIVOT/ITERATE decision.
G.
Documentation (Stages 16–19): Generates outline; writes draft; simulated peer review; revision.
H.
Finalization (Stages 20–23): Quality gate (Stage 20); knowledge archival; LaTeX export; 4-layer citation verification (arXiv ID, DOI, title match, relevance score).
Appendix E Optimization Trajectory Details
| Iter | Key Discovery | F1 | Type | |
| 0 | Naïve baseline | 0.117 | — | — |
| 1 | JSON response_format missing | 0.322 | +175% | Bug fix |
| 2 | BM25 hybrid | 0.464 | +44% | Architecture |
| 3 | Anti-hallucination prompting | 0.516 | +11% | Prompt |
| 4b | Evaluation format alignment | 0.543 | +5% | Format |
| 5 | MAU timestamp correction | 0.580 | +7% | Data repair |
| 6 | top-=30 + temporal hints | 0.577 | 0.5% | Hyperparam |
| 7b | Adaptive top- + metadata | 0.583 | +0.5% | Hyperparam |
| 8 | Forced exact-word copying | 0.551 | 5.5% | Reverted |
| 9 | Increased BM25 results | 0.575 | 1.4% | Reverted |
| benchmark validation | 0.598 | +411% | ||
| SimpleMem SOTA (Liu et al., 2026a) | 0.432 | |||
| Phase | Focus | F1 Range | Key Discovery | |
| 1 | Environment setup | 0.2540.353 | +39% | LLM upgrade + local embedding |
| 2 | Architecture | 0.3530.690 | +96% | Full-text retrieval + image BM25 |
| 3 | Fine-tuning | 0.6900.717 | +4% | Constraint position sensitivity |
| 4 | Scale validation | 0.7170.726 | +1% | Data completeness algorithms |
| 5 | Exact citation | 0.7260.771 | +6% | BM25 tokenization fix (+0.018) |
| 6 | Visual reasoning | 0.7710.789 | +2% | Image catalog + context |
| 7 | Plateau exploration | 0.7890.793 | +1% | Performance ceiling confirmed |
| Final (20 datasets) | 0.797 | +214% | ||
| MuRAG SOTA (Bei et al., 2026) | 0.697 | |||
Appendix F Full Iteration Logs
F.1 LoCoMo: Per-Category Performance Across Iterations
Table 8 shows how per-category F1 evolved across the optimization trajectory.
| Iter | Cat1 (MH) | Cat2 (T) | Cat3 (O) | Cat4 (SH) | Cat5 (A) |
| Baseline | 0.093 | 0.047 | 0.116 | 0.219 | 0.000 |
| Iter 1 | 0.206 | 0.117 | 0.229 | 0.381 | 0.447 |
| Iter 2 | 0.292 | 0.117 | 0.282 | 0.418 | 1.000 |
| Iter 3 | 0.351 | 0.335 | 0.346 | 0.444 | 1.000 |
| Iter 4b | 0.389 | 0.309 | 0.382 | 0.466 | 1.000 |
| Iter 5 | 0.398 | 0.487 | 0.388 | 0.445 | 1.000 |
| Iter 7b | 0.404 | 0.543 | 0.398 | 0.440 | 1.000 |
Notable patterns: Cat 5 (Adversarial) jumps from 0.447 to 1.000 after alignment with evaluation-standard refusal phrases in Iter 2. Cat 2 (Temporal) shows the most dramatic improvement (+0.496 total), driven primarily by the timestamp correction in Iter 5. Cat 3 (Open-ended) is our strongest category vs. prior SOTA, exceeding SimpleMem by +0.200.
F.2 Mem-Gallery: Detailed Phase-by-Phase Log
Phase 1: Environment Setup (Exp-000 to 001).
Starting from F1 = 0.254 with gpt-4.1-nano and CLIP-B/32 512-dim embeddings, the pipeline first upgraded the LLM to gpt-4o and switched to local all-MiniLM-L6-v2 embeddings (384d) after encountering API proxy errors. Result: F1 = 0.353 (+39%).
Phase 2: Architecture Breakthrough (Exp-002 to 004).
The largest single improvement (+53%) came from returning full original dialogue text instead of LLM summaries, which is counter-intuitive since summaries are traditionally preferred. The pipeline further added BM25 hybrid search (set-union merging) and created a dedicated image-caption BM25 index for visual question categories. Combined result: F1 = 0.690 (+96%).
Phase 3: Fine-Tuning (Exp-004b to 012).
The pipeline discovered that format constraint position (before vs. after the question) matters more than content at temperature=0. KR category improved +188% from repositioning alone. Several prompt-tuning experiments (Exp-008 to 011) failed and were reverted. Net result: F1 = 0.717 (+4%).
Phase 4: Scale Validation (Exp-014 to 018).
Validating on 14 datasets exposed data completeness issues (store_only not completed before qa_only). The pipeline learned that data completeness algorithmic optimization. Result: F1 = 0.726 (+1%).
Phase 5: Exact Citation (Exp-020 to 023).
Two key discoveries: (1) “Use the exact words and phrases from the conversation” instruction improved 12/15 datasets (+0.031 F1); (2) a simple BM25 tokenization fix (stripping punctuation: “sushi.” “sushi”) yielded +0.018 F1, which is more than 10 rounds of prompt engineering. Result: F1 = 0.771 (+6%).
Phase 6: Visual Reasoning Enhancement (Exp-026 to 027).
Augmenting the Image Catalog with dialogue context (full_text[:300]) for each image improved VR category by +0.087. Adding temporal ordering for TR/CD categories contributed another +0.006. Result: F1 = 0.789 (+2%).
Phase 7: Plateau Exploration (Exp-028 to 039b).
The pipeline explored model comparison (gpt-4.1 gpt-4o), per-category prompt refinements, and various BM25 configurations. After 4 independent runs yielding F1 in [0.791, 0.797], the pipeline correctly identified the performance ceiling ( due to random noise) and terminated. Final peak: F1 = 0.797.
Appendix G Prompt Catalog
This section documents representative prompts used throughout OmniMem, organized by system component. All prompts are reproduced from the source code; minor formatting adjustments are made for readability.
G.1 Core System Prompts
G.1.1 Answer Generation
The orchestrator uses a two-part prompt structure for answer generation: a system prompt that defines the assistant’s role, and a user prompt that injects retrieved context with structured output requirements. This design enforces JSON-mode output for reliable answer extraction.
The LLM is called with temperature=0.1 and response_format=json_object to ensure deterministic, parseable output. The structured JSON output was a key discovery of the autonomous pipeline: without explicit format enforcement, the model produced verbose natural-language answers that degraded F1 by 175% (see Iteration 1, Table 6).
G.1.2 Entity and Relation Extraction
The knowledge graph module extracts typed entities and directed relations from each MAU summary. The prompt enforces a structured JSON schema with confidence scores, enabling downstream entity resolution to filter low-confidence extractions.
The 7 entity types (Person, Location, Object, Event, Concept, Time, Organization) were selected to cover the dominant categories observed across both benchmarks. The multimodal guideline (“for visual descriptions, pay attention to spatial relationships”) enables the same extraction prompt to operate uniformly over text summaries generated from any input modality.
G.1.3 Query Intent Analysis
Before retrieval, the query processor optionally analyzes query intent to extract structured metadata. This analysis informs retrieval strategy selection (e.g., enabling temporal sorting for temporal queries, or activating visual search for image-related queries).
G.2 Modality Processor Prompts
OmniMem employs modality-specific prompts during ingestion to convert heterogeneous inputs into the unified MAU text summary. A key design principle is prompt minimalism: concise instructions yield shorter, more retrieval-friendly summaries at ingestion time, while detailed prompts are reserved for on-demand expansion during answer generation.
G.2.1 Image Captioning
Two levels of image captioning serve different stages of the pyramid retrieval pipeline:
The concise variant is used during ingestion (max_tokens=150), producing 10-token summaries stored in hot storage. The detailed variant is invoked only during Level 3 pyramid expansion (max_tokens=500), loading richer descriptions from cold storage on demand. This two-tier design reduces storage cost while preserving access to fine-grained visual detail when needed for complex queries.
G.2.2 Audio and Video Summarization
Audio content is first transcribed via speech-to-text, then summarized:
Video MAUs combine visual frame descriptions with audio transcripts into a unified summary:
Both prompts follow the minimalism principle: single-sentence summaries are sufficient for embedding-based retrieval in the first pyramid level, while the original transcripts and frame descriptions remain accessible in cold storage for deeper expansion.
G.3 Mem-Gallery Benchmark Prompts
Mem-Gallery evaluation employs a multi-level prompt architecture discovered through the autonomous optimization pipeline. A global system prompt defines the task context, category-specific format constraints control output formatting, and a dialogue agent prompt frames each question. The pipeline’s key discovery was that constraint positioning (before vs. after the question) significantly affects performance at low temperature, with per-category optimization yielding up to +188% improvement on individual categories (Section 3.3).
G.3.1 System Prompt
G.3.2 Category-Specific Format Constraints
Three of the nine categories receive explicit format constraints appended to the question. These constraints were iteratively refined by the pipeline to align model output with evaluation metrics:
The remaining six categories (FR, MR, TR, VR, TTL, KR) use the default system prompt without additional format constraints. The pipeline discovered that adding constraints to these categories either had no effect or degraded performance due to over-specification.
G.3.3 Dialogue Agent Prompt
Each question is framed by a dialogue agent prompt that is appended after the retrieved memory context:
G.3.4 Reasoning Model Adaptation
When using reasoning-class models (e.g., o1, o3) that tend to produce verbose chain-of-thought outputs, the pipeline appends a conciseness boost to the system prompt. This adaptation was discovered during Phase 7 when testing alternative backbones:
G.3.5 Memory Context Injection
Retrieved memories are formatted and injected into the prompt using the following template. For multimodal memories, image content is interleaved with textual context as base64-encoded content parts:

### 02 Apr 2026 v2

@@
-OmniMem: Autoresearch-Guided Discovery of Lifelong Multimodal Agent Memory
+Omni-SimpleMem: Autoresearch-Guided Discovery of Lifelong Multimodal Agent Memory
 Abstract
@@
 Building effective lifelong memory requires navigating a vast design space spanning architecture, retrieval strategies, prompt engineering, and data pipelines; this space is too large and interconnected for manual exploration or traditional AutoML to explore effectively.
-We deploy an autonomous research pipeline to discover OmniMem, a unified multimodal memory framework for lifelong AI agents.
+We deploy an autonomous research pipeline to discover Omni-SimpleMem, a unified multimodal memory framework for lifelong AI agents.
 Starting from a naïve baseline (F1 = 0.117 on LoCoMo), the pipeline autonomously executes experiments across two benchmarks, diagnosing failure modes, proposing architectural modifications, and repairing data pipeline bugs, all without human intervention in the inner loop.
-The resulting system achieves state-of-the-art on both benchmarks, improving F1 by +411% on LoCoMo (0.1170.598) and +214% on Mem-Gallery (0.2540.797) relative to the initial configurations.
+The resulting system achieves state-of-the-art on both benchmarks, improving F1 by +411% on LoCoMo and +214% on Mem-Gallery relative to the initial configurations.
 Critically, the most impactful discoveries are not hyperparameter adjustments: bug fixes (+175%), architectural changes (+44%), and prompt engineering (+188% on specific categories) each individually exceed the cumulative contribution of all hyperparameter tuning, demonstrating capabilities fundamentally beyond the reach of traditional AutoML.
 We provide a taxonomy of six discovery types and identify four properties that make multimodal memory particularly suited for autoresearch, offering guidance for applying autonomous research pipelines to other AI system domains.
-Code is available at this https://github.com/aiming-lab/OmniMem.
+Code is available at this https://github.com/aiming-lab/SimpleMem.
 1 Introduction
@@
 We ask whether this paradigm extends to complex, multi-component AI systems and answer affirmatively.
-We deploy AutoResearchClaw (Liu et al., 2026b), a 23-stage autonomous research pipeline, to discover OmniMem, a unified multimodal memory framework for lifelong AI agents.
+We deploy AutoResearchClaw (Liu et al., 2026b), a 23-stage autonomous research pipeline, to discover Omni-SimpleMem, a unified multimodal memory framework for lifelong AI agents.
 Starting from a naïve baseline (F1 = 0.117 on LoCoMo), the pipeline autonomously executes experiments across two benchmarks, iteratively diagnosing failure modes, proposing architectural modifications, repairing data pipeline bugs, and validating improvements, all without human intervention in the inner loop.
@@
 Critically, the most impactful discoveries are not hyperparameter adjustments: bug fixes (+175%), architectural changes (+44%), and prompt engineering (+188% on specific categories) each individually exceed the cumulative contribution of all hyperparameter tuning, demonstrating capabilities fundamentally beyond the reach of traditional AutoML.
-Among the pipeline’s most consequential discoveries are three architectural principles that define OmniMem.
+Among the pipeline’s most consequential discoveries are three architectural principles that define Omni-SimpleMem.
 First, selective ingestion: lightweight perceptual encoders measure the information novelty of each incoming signal and discard redundant content before storage, significantly reducing storage requirements.
@@
 Our key observation is that multimodal memory is particularly well-suited for autonomous research pipelines due to four properties: immediate scalar evaluation metrics enabling tight optimization loops, modular architecture allowing isolated component modification, fast iteration cycles (1–2 hours per experiment) supporting dozens of hypotheses within days, and version-controlled code modifications allowing failed experiments to be cleanly reverted.
-In summary, our primary contribution is OmniMem, a unified multimodal memory framework whose architecture and configuration are discovered through AutoResearchClaw and that achieves state-of-the-art results on both evaluated benchmarks.
+In summary, our primary contribution is Omni-SimpleMem, a unified multimodal memory framework whose architecture and configuration are discovered through AutoResearchClaw and that achieves state-of-the-art results on both evaluated benchmarks.
 Beyond the system itself, we provide a comprehensive taxonomy of autonomous discoveries across experiments, revealing that the highest-impact improvements lie beyond the reach of traditional AutoML, and we characterize the convergence behavior, failure modes, and automatic recovery patterns of the pipeline.
@@
 Our setting differs fundamentally: the “search space” includes not only hyperparameters and architectural choices, but also prompt engineering, data pipeline bug detection and repair, evaluation format alignment, and cross-component interaction diagnosis, all of which require natural language understanding and code modification capabilities beyond traditional AutoML.
-3 Autoresearch-Guided Discovery of OmniMem
+3 Autoresearch-Guided Discovery of Omni-SimpleMem
 In this section, we describe the autonomous optimization process and the system it produces.
-We first overview the pipeline (§3.1), then present the discovered OmniMem architecture (§3.2), followed by the benchmark-specific optimization strategy (§3.3).
+We first overview the pipeline (§3.1), then present the discovered Omni-SimpleMem architecture (§3.2), followed by the benchmark-specific optimization strategy (§3.3).
 3.1 Pipeline Overview
 As discussed in Section 1, the design space of multimodal memory systems is too large and interconnected for manual exploration to cover effectively.
-To address this, we deploy AutoResearchClaw (Liu et al., 2026b), a 23-stage autonomous research pipeline, to systematically optimize OmniMem.
+To address this, we deploy AutoResearchClaw (Liu et al., 2026b), a 23-stage autonomous research pipeline, to systematically optimize Omni-SimpleMem.
 The pipeline receives three inputs: (1) the SimpleMem (Liu et al., 2026a) codebase, a unimodal text-only lifelong memory framework, as a starting point, (2) two benchmark evaluation harnesses with quantitative metrics (F1), and (3) API access to LLM providers.
@@
 Once memories are ingested and stored as MAUs, the next challenge is how to retrieve them efficiently at query time.
-The second principle is progressive retrieval: rather than loading all retrieved content into the LLM context at once, OmniMem expands information in stages under explicit token budgets.
+The second principle is progressive retrieval: rather than loading all retrieved content into the LLM context at once, Omni-SimpleMem expands information in stages under explicit token budgets.
 Hybrid Dense-Sparse Search.
@@
 While hybrid search and pyramid retrieval handle queries that can be answered from individual MAUs, many real-world queries require reasoning over multiple connected facts (e.g., “What gift did I give to the person I met at the conference in March?”).
-The third principle is therefore structured knowledge: OmniMem maintains a knowledge graph that captures entities and relationships across all MAUs.
+The third principle is therefore structured knowledge: Omni-SimpleMem maintains a knowledge graph that captures entities and relationships across all MAUs.
 During MAU creation, an LLM extracts entities and directed relations from each summary, producing entity-relation triples.
@@
 4 Experiments
-We evaluate OmniMem along two dimensions: (1) the autonomous optimization process, specifically whether the pipeline discovers meaningful improvements across diverse benchmarks, and (2) final system quality, examining whether the discovered architecture achieves state-of-the-art results and whether individual components contribute meaningfully.
+We evaluate Omni-SimpleMem along two dimensions: (1) the autonomous optimization process, specifically whether the pipeline discovers meaningful improvements across diverse benchmarks, and (2) final system quality, examining whether the discovered architecture achieves state-of-the-art results and whether individual components contribute meaningfully.
 4.1 Experimental Setup
@@
 | SimpleMem | 0.318 | 0.195 | 0.235 | 0.308 | 0.802 | 0.432 | 0.535 | 0.348 | 0.310 | 0.468 | 0.390 | |
-| OmniMem | 0.556 | 0.365 | 0.255 | 0.641 | 0.835 | 0.598 | 0.797 | 0.449 | 0.366 | 0.627 | 0.505 | |
+| Omni-SimpleMem | 0.556 | 0.365 | 0.255 | 0.641 | 0.835 | 0.598 | 0.797 | 0.449 | 0.366 | 0.627 | 0.505 | |
 | GPT-4o-mini | MemVerse | 0.147 | 0.074 | 0.106 | 0.093 | 0.747 | 0.290 | 0.450 | 0.295 | 0.248 | 0.395 | 0.330 |
@@
 | SimpleMem | 0.300 | 0.128 | 0.178 | 0.312 | 0.891 | 0.404 | 0.498 | 0.318 | 0.290 | 0.435 | 0.368 | |
-| OmniMem | 0.544 | 0.196 | 0.177 | 0.588 | 0.779 | 0.519 | 0.749 | 0.403 | 0.334 | 0.583 | 0.465 | |
+| Omni-SimpleMem | 0.544 | 0.196 | 0.177 | 0.588 | 0.779 | 0.519 | 0.749 | 0.403 | 0.334 | 0.583 | 0.465 | |
 | GPT-4.1-nano | MemVerse | 0.146 | 0.061 | 0.169 | 0.115 | 0.711 | 0.256 | 0.470 | 0.308 | 0.255 | 0.410 | 0.340 |
@@
 | SimpleMem | 0.298 | 0.145 | 0.210 | 0.285 | 0.648 | 0.342 | 0.518 | 0.338 | 0.300 | 0.452 | 0.380 | |
-| OmniMem | 0.477 | 0.216 | 0.244 | 0.583 | 0.722 | 0.492 | 0.780 | 0.430 | 0.353 | 0.610 | 0.488 | |
+| Omni-SimpleMem | 0.477 | 0.216 | 0.244 | 0.583 | 0.722 | 0.492 | 0.780 | 0.430 | 0.353 | 0.610 | 0.488 | |
 | GPT-5.1 | MemVerse | 0.287 | 0.173 | 0.277 | 0.297 | 0.780 | 0.383 | 0.478 | 0.312 | 0.262 | 0.418 | 0.345 |
@@
 | SimpleMem | 0.305 | 0.178 | 0.272 | 0.305 | 0.807 | 0.418 | 0.538 | 0.350 | 0.312 | 0.470 | 0.395 | |
-| OmniMem | 0.598 | 0.367 | 0.307 | 0.676 | 0.747 | 0.613 | 0.810 | 0.460 | 0.374 | 0.639 | 0.515 | |
+| Omni-SimpleMem | 0.598 | 0.367 | 0.307 | 0.676 | 0.747 | 0.613 | 0.810 | 0.460 | 0.374 | 0.639 | 0.515 | |
 | GPT-5-nano | MemVerse | 0.208 | 0.203 | 0.168 | 0.252 | 0.741 | 0.366 | 0.478 | 0.315 | 0.262 | 0.420 | 0.345 |
@@
 | SimpleMem | 0.278 | 0.200 | 0.245 | 0.282 | 0.824 | 0.388 | 0.522 | 0.340 | 0.302 | 0.458 | 0.385 | |
-| OmniMem | 0.357 | 0.371 | 0.253 | 0.561 | 0.719 | 0.522 | 0.787 | 0.437 | 0.357 | 0.617 | 0.494 | |
+| Omni-SimpleMem | 0.357 | 0.371 | 0.253 | 0.561 | 0.719 | 0.522 | 0.787 | 0.437 | 0.357 | 0.617 | 0.494 | |
 Figure 3 visualizes the optimization trajectories.
@@
 4.3 Main Results
-To contextualize these results against existing memory systems, we conduct a controlled comparison of OmniMem against six baselines across five LLM backbones (GPT-4o, GPT-4o-mini, GPT-4.1-nano, GPT-5.1, and GPT-5-nano).
+To contextualize these results against existing memory systems, we conduct a controlled comparison of Omni-SimpleMem against six baselines across five LLM backbones (GPT-4o, GPT-4o-mini, GPT-4.1-nano, GPT-5.1, and GPT-5-nano).
 Table 1 reports per-category F1 on LoCoMo and five evaluation metrics on Mem-Gallery across all backbones.
-On LoCoMo, OmniMem achieves the highest overall F1 across all backbones, ranging from 0.492 (GPT-4.1-nano) to 0.613 (GPT-5.1), substantially outperforming SimpleMem (0.342–0.432), the current state-of-the-art on LoCoMo (Liu et al., 2026a).
-OmniMem dominates on multi-hop, single-hop, and open-domain categories, with particularly large margins on open-domain questions.
-On Mem-Gallery, OmniMem achieves F1 ranging from 0.749 to 0.810, consistently outperforming all memory baselines by a wide margin.
-SimpleMem is again the strongest baseline (F1 up to 0.538 with GPT-5.1), but still trails OmniMem by over 25 percentage points.
-These patterns confirm that OmniMem’s gains come from its architectural design (hybrid search, pyramid retrieval, knowledge graph augmentation) rather than from a single dominant component.
+On LoCoMo, Omni-SimpleMem achieves the highest overall F1 across all backbones, ranging from 0.492 (GPT-4.1-nano) to 0.613 (GPT-5.1), substantially outperforming SimpleMem (0.342–0.432), the current state-of-the-art on LoCoMo (Liu et al., 2026a).
+Omni-SimpleMem dominates on multi-hop, single-hop, and open-domain categories, with particularly large margins on open-domain questions.
+On Mem-Gallery, Omni-SimpleMem achieves F1 ranging from 0.749 to 0.810, consistently outperforming all memory baselines by a wide margin.
+SimpleMem is again the strongest baseline (F1 up to 0.538 with GPT-5.1), but still trails Omni-SimpleMem by over 25 percentage points.
+These patterns confirm that Omni-SimpleMem’s gains come from its architectural design (hybrid search, pyramid retrieval, knowledge graph augmentation) rather than from a single dominant component.
 4.4 Analysis
@@
 4.4.2 Efficiency
-OmniMem achieves 5.81 queries/sec with 8 parallel workers (3.5 faster than the fastest baseline), enabled by read-only FAISS and BM25 indices supporting concurrent lookup (Figure 4, Table 3).
-All baselines are bottlenecked by sequential LLM generation (85–97% of per-query time), while OmniMem parallelizes the retrieval-generation pipeline via thread-safe read-only indices.
+Omni-SimpleMem achieves 5.81 queries/sec with 8 parallel workers (3.5 faster than the fastest baseline), enabled by read-only FAISS and BM25 indices supporting concurrent lookup (Figure 4, Table 3).
+All baselines are bottlenecked by sequential LLM generation (85–97% of per-query time), while Omni-SimpleMem parallelizes the retrieval-generation pipeline via thread-safe read-only indices.
 | Method | q/s | Ret. | Gen. |
@@
 | Mem0 | 1.46 | 18 | 665 |
-| OmniMem (w=1) | 1.05 | 118 | 846 |
-| OmniMem (w=8) | 5.81 | 461 | 821 |
+| Omni-SimpleMem (w=1) | 1.05 | 118 | 846 |
+| Omni-SimpleMem (w=8) | 5.81 | 461 | 821 |
 | Ret./Gen. in milliseconds. | |||
 4.4.3 Case Study: Multi-Hop Retrieval
-We illustrate OmniMem’s retrieval pipeline on a real multi-hop query from LoCoMo that requires synthesizing facts across separate conversation sessions.
+We illustrate Omni-SimpleMem’s retrieval pipeline on a real multi-hop query from LoCoMo that requires synthesizing facts across separate conversation sessions.
 The query asks: “What subject have Caroline and Melanie both painted?” The gold answer is “sunsets,” but answering correctly requires retrieving each person’s painting history from different sessions and identifying the overlap.
 Hybrid search.
-Dense retrieval returns MAUs mentioning Caroline’s paintings (e.g., “Caroline painted a sunset scene”) and Melanie’s art projects (e.g., “Mel and her kids painted a sunset with a palm tree”), but these appear in separate sessions with different surrounding context.
+Dense retrieval returns MAUs mentioning Caroline’s paintings (e.g., “Caroline painted a sunset”) and Melanie’s art projects (e.g., “Mel and her kids painted a sunset with a tree”), but these appear in separate sessions with different surrounding context.
 BM25 recovers additional MAUs containing the keyword “paint” that rank lower in dense results.
@@
 5 Conclusion
-We have presented OmniMem, a unified multimodal memory framework whose architecture and configuration are discovered through AutoResearchClaw, an autonomous research pipeline.
-Starting from a naïve baseline (F1 = 0.117 on LoCoMo), the pipeline autonomously executed experiments in 72 hours, achieving state-of-the-art on both LoCoMo and Mem-Gallery.
-The highest-impact discoveries, including bug fixes (+175%), architectural changes (+44%), and prompt engineering (+188% on specific categories), require code comprehension and cross-component reasoning beyond the reach of traditional AutoML.
+We have presented Omni-SimpleMem, a unified multimodal memory framework whose architecture and configuration are discovered through AutoResearchClaw, an autonomous research pipeline.
+Starting from a naïve baseline, the pipeline autonomously executed experiments in 72 hours, achieving state-of-the-art on both LoCoMo and Mem-Gallery.
+The highest-impact discoveries, including bug fixes, architectural changes, and prompt engineering, require code comprehension and cross-component reasoning beyond the reach of traditional AutoML.
 Our taxonomy of six discovery types, together with the observation that multimodal memory is well-suited for autoresearch due to its scalar metrics, modular architecture, and fast iteration cycles, provides a roadmap for applying autonomous research pipelines to other complex AI system domains.
-Ethics Statement
-OmniMem persistently retains multimodal personal information (text, images, audio, video) over extended time horizons, raising important considerations regarding data privacy and consent.
-Real-world deployment necessitates explicit user consent, transparent retention policies, and robust mechanisms to review, modify, or delete stored memories.
-The knowledge graph component, which extracts and links entities across memory units, could enable profiling if deployed without adequate access controls and encryption.
-Regarding the autonomous research pipeline, all AutoResearchClaw experiments are conducted on established academic benchmarks with no real user data; deploying such pipelines on production systems would require additional oversight.
-We advocate for memory-augmented agents that prioritize user autonomy, data minimization, and the right to be forgotten.
-Reproducibility Statement
-We provide complete hyperparameters (Table 4), per-benchmark configurations (Table 5), and system architecture details in Appendix C.
-Full optimization trajectories with per-iteration metrics and decisions are documented in Appendix E.
-All prompts are reproduced verbatim in Appendix G, and the AutoResearchClaw pipeline stages are described in Appendix D.
-Both benchmarks, LoCoMo (Maharana et al., 2024) and Mem-Gallery (Bei et al., 2026), are publicly available; we evaluate under identical splits and protocols as prior work using standard metrics (token-level F1, exact match, BLEU).
-We open-source the complete framework, benchmark harnesses, and full experiment logs to enable independent replication.
 Appendix A Benchmark Details
@@
 It achieves strong performance on conversational tasks but is fundamentally text-only and bottlenecked by sequential LLM calls for each memory management decision.
-Appendix C OmniMem Technical Details
+Appendix C Omni-SimpleMem Technical Details
 C.1 System Architecture
-The complete OmniMem implementation comprises 13,300 lines of Python across 11 subpackages:
+The complete Omni-SimpleMem implementation comprises 13,300 lines of Python across 11 subpackages:
 Core (MAU data structure, event hierarchy, unified configuration)
@@
 Appendix G Prompt Catalog
-This section documents representative prompts used throughout OmniMem, organized by system component.
+This section documents representative prompts used throughout Omni-SimpleMem, organized by system component.
 All prompts are reproduced from the source code; minor formatting adjustments are made for readability.
@@
 G.2 Modality Processor Prompts
-OmniMem employs modality-specific prompts during ingestion to convert heterogeneous inputs into the unified MAU text summary.
+Omni-SimpleMem employs modality-specific prompts during ingestion to convert heterogeneous inputs into the unified MAU text summary.
 A key design principle is prompt minimalism: concise instructions yield shorter, more retrieval-friendly summaries at ingestion time, while detailed prompts are reserved for on-demand expansion during answer generation.